# Word-specific tonal realizations in Mandarin

### v1

@@ 2 Establishing word and meaning-specific pitch contours
This section describes how we addressed the first two predictions outlined in Section 1. We first modeled the pitch contours of spoken tokens of Mandarin disyllabic words with the RF tonal pattern, using generalized additive modeling. To explore Prediction 1, we evaluated the effectiveness of word type as a predictor of tonal realization, compared with the segment-related articulatory constraints previously described in the literature. To explore Prediction 2, we evaluated whether adding information about a word’s meaning in context would improve prediction of tonal realization, compared with word type alone. Finally, we triangulated the results of the generalized additive models with a random forest analysis. Sections 2.1 to 2.4 describe aspects of the methodology, Sections 2.5 to 2.7 report the results of the generalized additive models, Section 2.8 presents the random forest analysis, and Section 2.9 summarizes the Section 2 overall.
@@ 2.1 Generalized additive modeling
Classical analyses of pitch typically take measurements at various contour landmarks, such as maximum and minimum F0 values. However, since pitch actually varies continuously with time, such analyses miss much of the detail of the F0 contour. To better capture the complete shapes of tonal variations, we modeled the pitch contours of the tokens in our data using generalized additive models, henceforth GAMs (Wood, , 2017). GAMs relax the regression assumption that the relation between a predictor and response should be linear; instead, the model incorporates individual, potentially nonlinear relationships between each predictor variable and the response variable. For main effects, this relationship is estimated using functions known as smoothing splines (henceforth smooths), which can fit either a line or a (possibly wiggly) curve to the data, as required. Nonlinear interactions can be included using functions called tensor product smooths, which fit a wiggly (hyper)surface for the joint effect of two or more predictors. In addition, it is possible to include nonlinear random effects, for instance by using functions called factor smooths (Baayen et al., , 2022), which fit a wiggly curve for each level of the random factor, e.g. for each individual speaker in the case of a by-speaker factor smooth. Because GAMs model complex non-linear relationships, they make it possible to model F0 as a nonlinear function of time across an utterance, while also including other predictors known to affect pitch, such as speech rate and speaker gender. They can thus capture fine-grained modulations of pitch as time unfolds. For previous applications of GAMs to the analysis of pitch, see Kösling et al., (2013); Wieling et al., (2016); Chuang et al., (2021); Sun and Shih, (2021); Soskuthy, (2021). [^4] Examples of the application of GAMs to other phonetic data such as tongue movement trajectories obtained from electromagnetic articulography can be found in Wieling et al., (2016); Saito et al., (2023); Tomaschek et al., 2019b . For GAMs used for modeling chronometric data, see, e.g., Baayen et al., (2017, 2022); Heitmeier et al., 2023c .
@@
To illustrate our GAM-based modeling strategy, we created a toy dataset consisting of six disyllabic Mandarin word types with the RF tonal pattern. Their audio files were downloaded from Meng Dian, a publicly available Taiwanese online dictionary [^5] Available at https://racklin.github.io/moedict-desktop/addon.html.. In the audio files, a single speaker pronounces each word twice, so that we had a total of 12 tokens all from the same speaker. We measured the F0 of each token every 15 ms and then transformed these time points onto a normalized time scale of 0-1. The F0 values of one of the two renditions of each of the six words are plotted on this normalized time scale in the left-hand panel of Figure 1. It can be seen that the RF tonal sequence in Mandarin is realized with a small initial fall, followed by a rise, and finally a much larger fall. The dipping realization of T2 is consistent with previous findings that the rising portion of Mandarin T2 is usually preceded by a slight fall (Ho, , 1976; Tseng, , 1981; Shih, , 1988; Shen and Lin, , 1991; Xu, , 1997; Moore and Jongman, , 1997). [^6] The initial fall could also reflect dialectal variation specific to Taiwan Mandarin, as in this language, T2 is predominantly realized with a concave contour. This concave contour may have become a standardized realization that no longer reflects articulatory constraints (see e.g., Fon and Hsu, , 2007).
@@
Using the mgcv package (Wood, , 2017) for R (R Core Team, , 2022), we fitted a GAM to the toy dataset, with F0 as the dependent variable and normalized time as the only predictor. Including time as a predictor allows us to model the entire pitch contour of a tonal pattern by predicting F0 at regular intervals across a token of that pattern. The pitch contour predicted by the GAM, shown in the right-hand panel of Figure 1, captures the general trend with some precision, mirroring the raw data on the left. This graph is the model’s best estimate of the average population contour for words with the RF tonal pattern, given the 12 tokens in our toy dataset. However, the empirical contours show considerable variation around this average, even for a single speaker producing citation forms in isolation. This variability in realization is the focus of our study.
@@
To investigate whether individual tonal realizations might be related to the meaning of the carrier word token, we enriched the model with a by-word factor smooth, which is effectively a nonlinear, time-dependent, random effect for word type. In the present example, provided purely for illustration of the method, the by-word smooths are based on just two tokens of each word. [^7] For detailed discussion of the ways in which these smooths can be specified, and the accuracy of these smooths, see Baayen et al., (2022). This mixed model predicts, for each word type, a word-specific adjustment contour which has to be added to the population pitch contour to obtain the predicted pitch for a given word type. The by-word adjustment contours estimated by the GAM are visualized in the left-hand panel of Figure 2. The dotted line at $y=0$ is a reference line: an adjustment curve for a given word that followed this line would indicate that no adjustment is needed and that this word’s pitch is identical to the population contour. Deviations above this reference line indicate an upward F0 adjustment, and deviations below it indicate a downward adjustment. The word {CJK*}UTF8bsmi職業 zhíyè ‘profession’, for example, represented by a pink curve, requires an upward adjustment for the entire contour, although the amount of adjustment varies across time. When a given word’s adjustment contour is added to the general contour, we obtain its fitted contour, as shown in the right-hand panel of Figure 2. The dashed line in each graph represents the general contour which is, by definition, the same for all words. The red line (along with its confidence interval in blue) plots the fitted contour for the word in question. These fitted contours vary from word to word. For {CJK*}UTF8bsmi職業 zhíyè ‘profession’, presented in the right-most upper panel, the entire fitted contour is above the general average contour, as expected. The homophone pair {CJK*}UTF8bsmi程式 chéngshì ‘computer program’ and {CJK*}UTF8bsmi城市 chéngshì ‘city’, shown in the left and middle lower panels, have similar but not identical fitted contours, as would be expected if word meaning, as well as word form, plays a role in determining tonal realization.
@@
Figure 2: Toy dataset. The left-hand panel shows by-word adjustment contours from the toy model with only by-word factor smooth and normalized time as predictors. The right-hand panel plots the fitted contour for each word, with the predicted general contour (identical for all words) indicated by the dashed line.
@@ 2.2 Data
We used the Taiwan Mandarin spontaneous speech corpus (Fon, , 2004), which consists of about 30 hours of recorded interviews with 55 native Taiwan Mandarin speakers aged between 20 and 60 years, 31 identified as female and 24 identified as male. This corpus contains 94,783 disyllabic word tokens, 11,482 of which have the RF tonal pattern [^8] These tokens also include tone sandhi words containing {CJK*}UTF8bsmi一 and {CJK*}UTF8bsmi不 , which have T2 realizations when followed by T4 syllables.. These 11,482 tokens represent 707 orthographic word types. More than two-thirds of the tokens belong to one of five types: {CJK*}UTF8bsmi然後 ránhòu ‘and then’, {CJK*}UTF8bsmi時候 shíhòu ‘during which’, {CJK*}UTF8bsmi不會 búhuì ‘do not’, {CJK*}UTF8bsmi還是 háishì ‘still, or’ and {CJK*}UTF8bsmi一樣 yíyaǹg ‘likewise’. In order to avoid model predictions being heavily biased towards these five high-frequency words, we randomly sampled 300 tokens for each of these five types for inclusion in our dataset. [^9] The figure of 300 was an arbitrary choice at the upper end of the frequency range for the other words in the dataset. We furthermore excluded words with fewer than 20 occurrences from the dataset, in order to avoid overfitting to low-frequency words. As a consequence, our initial dataset comprised 9,496 tokens across 53 word types.
@@
Subsequently, we extracted the sound files of these tokens and measured their F0 values using Praat (Boersma and Weenink, , 2019). F0 values were taken every 15 ms, and therefore tokens of longer duration had more measurement datapoints than shorter tokens. In order to make sure every token had sufficient datapoints for model fitting, we excluded extremely short tokens with fewer than six datapoints, which constituted about 5% of the data. Next, we removed tokens where an F0 extraction error was likely. F0 extraction errors usually result from pitch halving or doubling, and lead to abrupt big changes in the recorded F0 values. We therefore first obtained, for each token, all the F0 differences between consecutive measurements, and then calculated the standard deviation of these difference values. The standard deviation is large when F0 measurements are discontinuous and fluctuate abruptly. Tokens with standard deviations greater than the 9th decile of the distribution were considered to be outliers, hence likely to involve extraction errors, and were excluded from further analyses. Finally, two words were excluded because their tokens were all contributed by only one speaker.
@@
The final dataset for the first analysis, reported below in Section 2.6, consisted of a total of 3,778 tokens representing 51 word types. Since these types do not include any heterographic homophones, there is a one-to-one correspondence between the orthographic labels of the tokens in our data and their canonical spoken forms. We therefore assume that tokens with the same label bear some similarity to one another in both form and meaning, i.e. belong to the same word type. [^10] Heterographic homophones are relatively rare amongst Mandarin disyllabic words. By definition, the spoken forms of disyllabic words have more possible combinations of segments and canonical tones than monosyllabic words do; this means there is less need to reuse the same canonical spoken forms for different meanings.
@@ 2.3.1 Core predictors
Normalized time: The points in time at which F0 measurements were taken (at 15 ms intervals) were, for each token, transformed into a normalized time scale of 0-1 to produce the variable time.
@@ 2.3.3 Context-related controls
Speech rate: The shape of tonal contours is modulated by how fast speakers talk. Generally, the faster the speech rate, the more likely it is that tonal contours will deviate from their underlying shapes (Cheng and Xu, , 2015; Tang and Li, , 2020). We therefore calculated speech_rate for each token, defined as syllables per second over the time period from four words before to four words after the target token, inclusive.
@@
Utterance position: The realization of tone in an utterance is also influenced by sentence intonation (Ho, , 1976; Tseng, , 1981; Shen, , 1989; Shen, 1990a, ). For example, statement intonation is often characterized by a downward trend, resulting in pitch declination (Shih, , 1997). Question intonation, on the other hand, can potentially lead to a final rise, although this largely depends on the syntactic structure and/or emotive force of the question concerned (Lee, , 2005; Chuang et al., , 2007). For the current study, we simply calculated the normalized position of each token in the relevant utterance. We defined an utterance as a sequence of words preceded and followed by a perceivable pause (regardless of duration), as indicated by the labels provided in the corpus. The variable utterance_position is the position at which a given token occurs in an utterance divided by the total number of words in that utterance. This predictor is therefore bounded between 0 and 1. For utterances with only one word, the utterance position was set to 1.
@@
where Freq denotes word frequency in the corpus of Taiwan Mandarin (Fon, , 2004).

### v2

@@ 2 Establishing word and meaning-specific pitch contours
This section describes how we addressed the first two predictions outlined in Section 1. We modeled the pitch contours of spoken tokens of Mandarin disyllabic words with the RF tonal pattern, using generalized additive modeling. To explore Prediction 1, we evaluated the effectiveness of word type as a predictor of tonal realization, compared with the segment-related articulatory constraints previously described in the literature. To explore Prediction 2, we evaluated whether adding information about a word’s meaning in context would improve prediction of tonal realization, compared with prediction based on word type alone. Sections 2.1 to 2.4 describe aspects of the methodology, Sections 2.5 to 2.7 report the results of the generalized additive models, and Section 2.8 summarizes Section 2 overall.
@@ 2.1 Generalized additive modeling
Classical analyses of pitch typically take measurements at various contour landmarks, such as maximum and minimum F0 values. However, since pitch actually varies continuously with time, such analyses miss much of the detail of the F0 contour. To better capture the complete shapes of tonal variations, we modeled the pitch contours of the tokens in our data using generalized additive models, henceforth GAMs (Wood,, 2017). GAMs relax the regression assumption that the relation between a predictor and response should be linear; instead, the model incorporates individual, potentially nonlinear relationships between each predictor variable and the response variable. For main effects, this relationship is estimated using functions known as smoothing splines (henceforth smooths), which can fit either a line or a (possibly wiggly) curve to the data, as required. Nonlinear interactions can be included using functions called tensor product smooths, which fit a wiggly (hyper)surface for the joint effect of two or more predictors. In addition, it is possible to include nonlinear random effects, for instance by using functions called factor smooths (Baayen et al.,, 2022), which fit a wiggly curve for each level of the random factor, e.g. for each individual speaker in the case of a by-speaker factor smooth. Because GAMs model complex non-linear relationships, they make it possible to model F0 as a nonlinear function of time across an utterance, while also including other predictors known to affect pitch, such as speech rate and speaker gender. They can thus capture fine-grained modulations of pitch as time unfolds.
@@
To illustrate our GAM-based modeling strategy, we created a toy dataset consisting of six disyllabic Mandarin word types with the RF tonal pattern. Their audio files were downloaded from Meng Dian, a publicly available Taiwanese online dictionary. In the audio files, a single female speaker pronounces each word twice, so that we had a total of 12 tokens, all from the same speaker. We used the To Pitch (cc) command in Praat (Boersma and Weenink,, 2019) to estimate F0 values for each token. Because the speaker is female, we set the pitch floor at 50 Hz and the pitch ceiling at 400 Hz. To optimize the F0 estimation, we set the time step at 0.001 seconds and used the most accurate method available, namely a Gaussian window; other parameters were left at the default values. This gave us F0 values at one millisecond intervals for the voiced sections of the tokens. Next, we used Praat’s To PointProcess command to obtain the time points of the glottal pulses in the voiced sections of the tokens. We then extracted the F0 values corresponding to the time points of the glottal pulses. Since the time points of glottal pulses do not necessarily correspond exactly to the one-millisecond measurement intervals, we used linear interpolation between adjacent measurements to estimate F0 for the glottal pulses. At this stage no F0 values were interpolated for the voiceless sections of the tokens. Because the tokens varied in duration, we transformed the time points of the measurements onto a normalized time scale of 0-1.555 Time normalization enables us to focus on the shape of pitch contours. In order to take into account possible effects of word duration on the realization of tone, in the more comprehensive analyses reported below, token duration is included as a covariate, in interaction with normalized time. Note that by reversing the time normalization, the predicted curves can straightforwardly be back-transformed to the original time scale. The F0 values of one of the two renditions of each of the six words are plotted on this normalized time scale in the left-hand panel of Figure 1. Although there are no data points for the voiceless segments, it can be seen that the RF tonal sequence in Mandarin is realized with a small initial fall, followed by a rise, and finally a much larger fall. The dipping realization of T2 is consistent with previous findings for laboratory speech that the rising portion of Mandarin T2 is usually preceded by a slight fall (Ho,, 1976; Tseng,, 1981; Shih,, 1988; Shen and Lin,, 1991; Xu,, 1997; Moore and Jongman,, 1997). [^6] The initial fall could also reflect dialectal variation specific to Taiwan Mandarin, as in this language, T2 is predominantly realized with a concave contour. This concave contour may have become a standardized realization that no longer reflects articulatory constraints (see e.g., Fon and Hsu,, 2007).
Using the mgcv package (Wood,, 2017) for R (R Core Team,, 2022), we fitted a GAM to the toy dataset, with F0 as the dependent variable and normalized time as the only predictor. Including time as a predictor allows us to model the entire pitch contour of a tonal pattern by predicting F0 across the whole timespan of a token of that pattern. The GAM predicts F0 values not only for voiced sections but also for voiceless parts of the tokens by estimating what the pitch contour would likely be if the voiceless segments were voiced. In adopting this modeling strategy, our goal is not curve fitting, but cognitive modeling of pitch contours. We theorize that a speaker producing a word has a pitch contour for the whole word or even the whole utterance in mind. We further assume that the cognitive planning underlying the production of pitch contours is continuous, rather than a step function with jumps to zero Hz for voiceless segments. From our theoretical perspective, it is only at the stage of articulation that voiceless segments mask the internally projected pitch development.
@@
The pitch contour predicted by the GAM, shown in the right-hand panel of Figure 1, captures the general trend with some precision, mirroring the raw data on the left. This graph is the model’s best estimate of the average population contour for words with the RF tonal pattern, given the 12 tokens in our toy dataset. However, the empirical contours show considerable variation around this average, even for a single speaker producing citation forms in isolation. This variability in realization is the focus of our study.
@@
To investigate whether individual tonal realizations might be related to the meaning of the carrier word token, we enriched the model with a by-word factor smooth, which is effectively a nonlinear, time-dependent, random effect for word type. In the present example, provided purely for illustration of the method, the by-word smooths are based on just two tokens of each word. [^7] For detailed discussion of the ways in which these smooths can be specified, and the accuracy of these smooths, see Baayen et al., (2022). This mixed model predicts, for each word type, a word-specific adjustment contour which has to be added to the population pitch contour to obtain the predicted pitch for a given word type.
The by-word adjustment contours estimated by the GAM are visualized in the left-hand panel of Figure 2. The dotted line at $y=0$ is a reference line: an adjustment curve for a given word that followed this line would indicate that no adjustment is needed and that this word’s pitch is identical to the population contour. Deviations above this reference line indicate an upward F0 adjustment, and deviations below it indicate a downward adjustment. The word {CJK*}UTF8bsmi職業 zhi2ye4 ‘profession’, for example, represented by a green curve, requires an upward adjustment for the entire contour, although the amount of adjustment varies across time. When a given word’s adjustment contour is added to the general contour, we obtain its fitted contour, as shown in the right-hand panel of Figure 2. The dashed line in each graph represents the general contour which is, by definition, the same for all words. The red line (along with its confidence interval in blue) plots the fitted contour for the word in question. These fitted contours vary from word to word. For {CJK*}UTF8bsmi職業 zhi2ye4 ‘profession’, presented in the right-most upper panel, the entire fitted contour is above the general average contour, as expected. The homophone pair {CJK*}UTF8bsmi城市 cheng2shi4 ‘city’ and {CJK*}UTF8bsmi程式 cheng2shi4 ‘computer program’, shown in the left and middle lower panels, have similar but not identical fitted contours, as would be expected if word meaning, as well as word form, plays a role in determining tonal realization.
Because the GAM receives no input for voiceless sections of the tokens, where there is no actual pitch, the model’s estimations for these sections are more uncertain than for the voiced sections. Similarly, word-specific partial effects for voiceless sections tend towards zero, i.e., no effect, again because there is no data. For instance, in Figure 2, the panel for {CJK*}UTF8bsmi程式 cheng2shi4 ‘program’ shows that the GAM produces a wide confidence interval for the two voiceless consonants, exactly as it should. Because there is no data for the voiceless segments of a given token, the model falls back on the overall smooth of time in these intervals, i.e., the general rise-fall tonal pattern. Hence, in Figure 2, it can be seen that the wide confidence intervals for voiceless sections of the tokens tend towards including the dotted line that represents the average contour.
@@
Figure 2: The left-hand panel shows by-word adjustment contours from the toy model with only by-word factor smooth and normalized time as predictors. The right-hand panel plots the fitted contour for each word, with the predicted general contour (identical for all words) indicated by the dashed line.
@@ 2.2 Data
We used the Taiwan Mandarin spontaneous speech corpus, which is one of a pair of corpora created by researchers at the National Taiwan University (see Fon,, 2004, for a description of the general method in the context of the Southern Min corpus). The Mandarin corpus consists of about 30 hours of recorded interviews with 55 native Taiwan Mandarin speakers aged between 20 and 60 years, 31 self-identified as female and 24 self-identified as male, recruited through snowball sampling starting with friends, acquaintances or relatives of the team. The recordings were made over several years between 2000 and 2010, using high-quality microphones and digital recorders in quiet locations, whenever possible in a soundproof laboratory. Before the interview, participants were told that the purpose was to understand their views on changes in various aspects of life. The interviewers aimed not to engage in conversation with participants but to elicit longer monologues about their personal experiences in childhood, school, work, relationships and elsewhere. After the interview, participants were debriefed and asked for permission to include their recording in the corpus. If permission was granted, the researchers orthographically transcribed the recording using Chinese characters. The word segmentation system developed by Academia Sinica (Ma and Chen,, 2003) was used to identify word boundaries in the orthographic representation, and dictionaries were used to add information about the canonical tones of the words identified. The character transcriptions were then romanized in order to run a forced aligner (Easyalign: Goldman,, 2011) to match the transcription to syllable boundaries in the audio files. The resulting alignment was manually checked and where necessary corrected.
@@
The corpus described in the previous paragraph contains 94,783 disyllabic word tokens, 11,482 of which have the RF tonal pattern [^8] These tokens also include tone sandhi words containing {CJK*}UTF8bsmi一 yi1 and {CJK*}UTF8bsmi不 bu4, which have T2 realizations when followed by T4 syllables.. These 11,482 tokens represent 707 orthographic word types. However, more than two-thirds of the tokens belong to one of five types: {CJK*}UTF8bsmi然後 ran2hou4 ‘and then’, {CJK*}UTF8bsmi時候 shi2hou4 ‘during which’, {CJK*}UTF8bsmi不會 bu2hui4 ‘do not’, {CJK*}UTF8bsmi還是 hai2shi4 ‘still, or’ and {CJK*}UTF8bsmi一樣 yi2yang4 ‘likewise’. In order to avoid model predictions being heavily biased towards these five high-frequency words, we randomly sampled 300 tokens for each of these five types for inclusion in our dataset. [^9] The figure of 300 was an arbitrary choice at the upper end of the frequency range for the other words in the dataset. We also excluded types with fewer than 20 occurrences from the dataset, in order to avoid overfitting to low-frequency words. As a consequence, our initial dataset comprised 4,516 tokens across 53 word types.
@@
We extracted the sound files for our 4,516 tokens and measured their F0 values using the method described in Section 2.1 for the toy dataset. For speakers identified as male, we set the pitch floor and ceiling to 50 and 400 Hz respectively. In order to make sure every token had sufficient data points for model fitting, we excluded extremely short tokens with fewer than six data points, which constituted about 5% of the data. Next, we removed tokens where an F0 extraction error was likely. F0 extraction errors usually result from pitch halving or doubling, and lead to abrupt big changes in the recorded F0 values. We therefore first obtained, for each token, all the F0 differences between consecutive measurements, and then calculated the standard deviation of these difference values. The standard deviation is large when F0 measurements are discontinuous and fluctuate abruptly. Tokens with standard deviations greater than the 9th decile of the distribution were considered to be outliers, hence likely to involve extraction errors, and were excluded from further analyses. Finally, two words were excluded because their tokens were all contributed by only one speaker.
@@
The final dataset for the first analysis, reported below in Section 2.6, consists of a total of 3,778 tokens representing 51 word types. Since these types do not include any heterographic homophones, there is a one-to-one correspondence between the orthographic labels of the tokens in our data and their canonical spoken forms. We therefore assume that tokens with the same label bear some similarity to one another in both form and meaning, i.e. belong to the same word type. However, because these tokens were extracted from spontaneous speech, not every speaker produced every type. To ensure that no word was completely nested under speaker or vice versa, we checked that every type was produced by multiple speakers and that every speaker produced multiple types. In the dataset of 3,778 tokens, the median number of speakers per word type is 20, the mean is 24.45, and the range is 5–52. The median number of word types per speaker is 23, the mean is 22.67, and the range is 12–35.
In the dataset described above, the F0 values are positively skewed. We therefore log-transformed them to create a response variable with a distribution closer to Gaussian, a requirement for modeling with Gaussian GAMs. [^10] For studies of auditory comprehension, several psychoacoustic scales are available, such as the Bark scale, which directly relate to human perception of pitch. However, as the interest of the present study is primarily in the pitch that speakers produced, we wanted to stay as close as possible to the physical signal and therefore simply log transformed the Hertz values. Note that despite voiceless gaps in the data for certain tokens, the overall distribution of data points from all tokens combined is dense right across the [0,1] interval of normalized time. The model is therefore able to make accurate predictions for all tokens within this range, including those with voiceless segments, albeit with variations in confidence as described for the toy dataset in Section 2.1.
@@ 2.3.1 Core predictors
Normalized time: The points in time at which F0 measurements were taken were, for each token, transformed into a normalized time scale of 0-1 to produce the variable time.
@@ 2.3.3 Context-related controls
Duration: The shapes of tonal contours are influenced by the time available to articulate them. Under time pressure, i.e., when words are spoken quickly, speakers do not have enough time to realize the full tonal contours, which are therefore more likely to deviate from their underlying shapes (Cheng and Xu,, 2015; Tang and Li,, 2020). We measured the duration of each token in seconds. As the distribution of these measurements was heavily skewed to the right, we log-transformed them before conducting GAM analyses. In our models, the variable duration is the log-transformed token duration. [^12] In our data, token duration is strongly correlated with speech rate (r =-0.57, p ¡ 0.0001). In other words, these two variables could not meaningfully be included as predictors in the same model (cf. Section 2.4). Since duration turned out to predict F0 better than speech rate did, we report models with duration.
@@
Utterance position: The realization of tone in an utterance is also influenced by sentence intonation (Ho,, 1976; Tseng,, 1981; Shen,, 1989; Shen, 1990a, ). For example, statement intonation is often characterized by a downward trend, resulting in pitch declination (Shih,, 1997). Question intonation, on the other hand, can potentially lead to a final rise, although this largely depends on the syntactic structure and/or emotive force of the question concerned (Lee,, 2005; Chuang et al.,, 2007). For the current study, we simply calculated the normalized position of each token in the relevant utterance. We defined an utterance as a sequence of words preceded and followed by a perceivable pause (regardless of duration), as indicated by the labels provided in the corpus. The variable utterance_position is the position at which a given token occurs in an utterance divided by the total number of words in that utterance. This predictor is therefore bounded between 0 and 1. For utterances with only one word, the utterance position was set to 1.131313In many languages, F0 at the start of an utterance is related to the length of the utterance, with higher initial F0 for longer utterances. We therefore tried including utterance length in our models. Although overall model fit improved, the concurvity score of utterance length was as high as 0.84, suggesting that a large portion of the effect can be explained by other factors in the model (cf. Section 2.6.1). In particular, utterance length is negatively correlated with token duration (r =-0.23), so keeping both variables would have rendered their effects uninterpretable. We chose to remove utterance length and keep duration.
@@
where $P(w_{n}|w_{n-1})$ is the probability of a word occurring given the previous word, $P(w_{n}|w_{n+1})$ is the probability of a word occurring given the following word, and Freq denotes word frequency in the corpus of Taiwan Mandarin.